# Incubulable hyperbolic $3$-pseudomanifold groups

### v1

@@ Abstract.
We construct compact hyperbolic $3$-manifolds with totally geodesic boundary, such that the closed $3$-pseudomanifolds obtained by coning off the boundary components are negatively curved and contain locally convex subspaces whose fundamental groups have property (T). In particular, the fundamental groups of these $3$-pseudomanifolds are word hyperbolic but not cubulable.
@@ 1. Introduction
To say a group is cubulable is to say it admits a proper and cocompact action by isometries on some $\textrm{CAT}(0)$ cube complex. Every cocompact Kleinian group is cubulable by [BW12, Theorem 1.5]. And a cubulable word hyperbolic group with $2$-sphere boundary is virtually cocompact Kleinian by [Ha15, Theorem 1.10] (generalizing [Mar13]).
We do not directly speak to Question 1.1, but we study the corresponding question for groups with Pontryagin sphere boundary. The Pontryagin sphere is a certain inverse limit of surfaces which appears as the boundary at infinity of many $\textrm{CAT}(0)$ 33-dimensional complexes (universal covers of $3$-pseudomanifolds, see below). It is like a sphere in that it is a connected $2$-dimensional compactum whose $2$-dimensional Čech cohomology is cyclic. However its $1$-dimensional cohomology is infinitely generated. We refer the reader to [JA91, DR99, FI03, SW20].
@@
Classical examples of word hyperbolic groups with Pontryagin sphere boundary occur for instance as finite index subgroups of a right-angled Coxeter group (RACG) whose nerve is a flag-no-square triangulation of a connected closed orientable surface, and via Charney–Davis strict hyperbolization [DA08, CD95]. Both sources of examples yield groups that are cubulable (see [DA08, LR24]). Nonetheless, we answer Question 1.2 in the negative; see (1) in Corollary 1.4.
@@
The existence of an infinite subgroup with property (T) is a well-known obstruction to the existence of nice geometric actions, and makes these groups quite different from $3$-manifold groups. More precisely, we have the following. (See [BHV08] for the definitions of the Haagerup and (T) properties, and [AG08] for the definition of RFRS.)
@@
The non-cubulability of $\pi_{1}(\widehat{M}_{n})$ does not automatically imply the non-cubulability of fundamental groups of $3$-pseudomanifolds obtained by coning off the boundary components of an arbitrary finite-sheeted cover of $M_{n}$. Indeed, one may use Wise’s Malnormal Special Quotient Theorem [Wise21, Theorem 12.2] together with the cubulability of $\pi_{1}(M_{n})$ [Wise21, Theorem 17.14] to see that each $M_{n}$ has some finite sheeted cover $M_{n}^{\prime}$ so that $\pi_{1}(\widehat{M}_{n}^{\prime})$ is cubulable.
@@
The Gromov boundary of the groups $\pi_{1}(\widehat{M}_{n})$ in Theorem 1.3 is the tree of manifolds defined by closed orientable surfaces of positive genus, i.e., a Pontryagin sphere, see [KM, Theorem A.1] and [SW20]. Moreover, the Gromov boundary of the infinite quasiconvex subgroup with property (T) has limit set a Menger curve, see [KK00].
@@
The groups $\pi_{1}(\widehat{M}_{n})$ in Theorem 1.3 split over quasiconvex surface subgroups (see Proposition 5.1), so they do not have property (T). In particular, they are $3$-dimensional word hyperbolic groups without (T) and without Haagerup, whose boundary is connected and has cyclic top-dimensional Čech cohomology (since the boundary is a Pontryagin sphere, see [SW20, Theorem 6.2]). See [LR25, Remark 5.6] for other examples in dimension $\geq 9$, whose boundary is a topological sphere. For examples of non-Kleinian hyperbolic 3-pseudomanifold groups that have the Haagerup property, and even act convex cocompactly on $\mathbb{H}^{n}$ for some $n>3$, see the RACGs considered in [DLMR25].
@@
In §2 we fix our notation and terminology about the construction of pseudomanifolds as cone-offs of manifolds and about orbifolds. Moreover, we collect some preliminary material from [LMW19] about the $2$-dimensional simplicial complexes $T_{n}$ whose fundamental groups have property (T). In §3 we present the construction of a particular hyperbolic $3$-orbifold with boundary, and in §4 we show how to construct suitable finite covers for which we can obtain quantitative control on various geometric quantities. The orbifold structure constructed in §3 is given by orthogonal mirrors. This is reminiscent of the Davis reflection trick [DA83]. However, we use the mirror structure to construct manifolds with boundary instead of closed manifolds. The geometric control enables us to use the results from [KM] to construct negatively curved metrics on the cone-offs of such covers. The proofs of the main theorem and corollary stated in the Introduction are presented in §5. We end the paper with some open questions in §6.
Acknowledgments. J.M. was partially supported by the Simons Foundation, grant #942496. L.R. was partially supported by INDAM-GNSAGA. We thank Chris Hruska for pointing us to useful references.
@@ 2.1. Coning off
Here we state the results from [KM] which allow us to put negatively curved (i.e., locally $\textrm{CAT}(k)$ for some $k<0$) metrics on our $3$-pseudomanifolds. Let $M$ be a compact $3$-manifold with boundary. The cone-off is the space
@@
Sufficient geometric conditions are given in [KM] for $\widehat{M}$ to admit a negatively curved metric $\hat{d}$ and for the induced cone-off of a subspace $Z$ to be locally convex. To state the conditions we need a definition from [KM]:
@@
[KM, Theorem A] Let $M$ be a compact hyperbolic manifold with totally geodesic boundary, let $b$ be a positive number less than $\operatorname{BW}_{M}(\partial M)$ and let $c>\pi/\sinh(b)$. Suppose
@@
The following is a special case of a result proved in [KM]:
@@
[KM, Theorem B] Let $M$, $b$, and $c$ satisfy the hypotheses of Theorem 2.2 and let $\hat{d}$ be the negatively curved metric on $\widehat{M}$ constructed in that theorem. Suppose $Z\subset M$ is closed, locally convex, is a tame product near $N_{b}(\partial M)$, and that
@@ 2.2. Orbifolds
For an introduction to the general notion of a smooth orbifold see [BB12, BH99, BMP, CHK, Kap01]. We briefly recall the idea: An $n$-orbifold (possibly with boundary) consists of a topological space (the underlying space) together with an open cover by sets of the form $U=\tilde{U}/\Gamma$, where $\tilde{U}$ is an open subset of $\mathbb{R}^{n}$ or $\mathbb{R}^{n-1}\times[0,\infty)$, and $\Gamma$ is a finite group of diffeomorphisms with nonempty fixed point set. (Smoothness also implies that a nonempty intersection between these sets gives rise to smooth transition maps.) If a point $x$ in the underlying space is the image of a fixed point of $\Gamma$, we say that the isotropy group of $x$ is $\Gamma$. (Strictly speaking we replace $\Gamma$ by the finite subgroup of $O(n)$ given by taking the derivative of its elements; this gives a well-defined group up to conjugacy in $O(n)$.) The boundary $\partial X$ is the (n1)(n-1)-orbifold whose underlying space consists of those points where the orbifold chart always comes from $\mathbb{R}^{n-1}\times[0,\infty)$, and whose isotropy groups are the isotropy groups for $X$, restricted to $\mathbb{R}^{n-1}\times\{0\}$. The isotropy groups of the points in the boundary lie in O(n1)O(n-1). The orbifold $X$ is said to be closed if its underlying space is compact and $\partial X$ is empty.
@@ 2.3. The seed with property (T)
We describe here a construction which is closely related to the examples in [LMW19]. In particular we will describe some fundamental groups of triangle complexes which are subgroups of finite index in some of the groups considered there.
@@
Let $\Theta$ be the $1$-complex with two vertices and $k$ edges between them. The following is a consequence of work of Lubotzky–Phillips–Sarnak [LPS88] and Margulis [MA88].
@@
This is explained in [LMW19], following closely [Lub94, Chapter 7]. For the fact that the graphs can be taken in a tower, see the bottom of page 64 of [LMW19]. The reason is that the subgroups of $\pi_{1}(\Theta)$ corresponding to the covers $\Lambda_{n}\to\Theta$ arise as congruence subgroups $\Gamma(2q_{n})$ for an arithmetic group $\Gamma\cong\pi_{1}(\Theta)$, and the numbers $q_{n}$ can be taken to be successive powers of a prime $q$ where k1k-1 is not a square modulo $q$. (For example if k1=17k-1=17 we could take $q_{n}=3^{n}$ or $q_{n}=5^{n}$.) ∎
@@
Since $\tau$ is induced by a graph automorphism of $\Theta$, all the vertex links in $\widehat{T}_{n}$ are isomorphic as graphs to the cover $\Lambda_{n}$. By Theorem 2.6, for $n$ large enough we can assume that the link of every vertex in $\widehat{T}_{n}$ has girth at least $6$. Now, if we identify each triangle of $\widehat{T}_{n}$ with an equilateral Euclidean triangle, then the induced metric is locally $\textrm{CAT}(0)$, so $\widehat{T}_{n}$ is aspherical. This asphericity, together with the condition on the Laplacian of $\Lambda_{n}$ in Theorem 2.6, allows us to apply [BS97, Corollary 1] (or [Zuk96, Théorème 1]) to deduce that $\pi_{1}(\widehat{T}_{n})$ has property (T). ∎
@@ 3. The handlebody $H_{0}$ and its hyperbolic orbifold structure
To recall the strategy of our main construction: We will first take the cover $p:H_{n}\to H_{0}$ corresponding to a cover $T_{n}\to T_{0}$ from Definition 2.7, with $n$ large enough to apply Lemma 2.8. We will then take an orbifold cover $M$ of $H=H_{n}$ to which $H=H_{n}$ lifts homeomorphically, and then cone off $\partial M$. Choosing these covers carefully, and using the results from [KM] (see §2.1) the inclusion of the induced cone-off $\widehat{H}$ into $\widehat{M}$ will have $\pi_{1}$-image isomorphic to $\pi_{1}(\widehat{T}_{n})$, which has property (T) by Lemma 2.8.
@@ 3.1. The hyperbolic orbifold structure on $H_{0}$
The underlying space of $B$ is a closed ball, but the orbifold structure suggests an abstract polyhedron, with dihedral angles of $\pi/2$ at all edges involving the unmirrored face $F_{B}$, and $\pi/k$ at each edge with isotropy group dihedral of order $2k$. One can see by direct inspection that this polyhedron satisfies the hypotheses of Andreev’s Theorem (see [RHD07, Theorem 1.4], which corrects some mistakes in the original [And70]). In particular, there is a unique hyperbolic polyhedron with the specified dihedral angles, and we identify this polyhedron with $B$. Mirroring all the faces with an acute adjacent dihedral angle gives a hyperbolic orbifold with the same orbifold structure specified above on $B$. In other words the polyhedron determines a hyperbolic structure on the orbifold $B$. Since the dihedral angles adjacent to the boundary face $F_{B}$ are $\pi/2$, this structure has totally geodesic boundary.
@@ 4.1. The orbifold $H=H_{n}$ covering $H_{0}$
Each component of $\partial H_{n}$ is a $2$-dimensional thickening of a cover of a theta graph. A cover of a theta-graph can be non-planar (eg $K_{3,3}$), so it is possible that these components have positive genus. Indeed, they must: if they were planar, then the coned-off complex would be homotopy equivalent to a $3$-manifold with boundary, so its fundamental group could not contain an infinite subgroup with property (T), see [FU99].
@@ 4.3. The manifold $M$ orbifold covering $H$
We want to construct a manifold $M$ (with empty orbifold locus) that is an orbifold cover of the orbifold $H$, so that the underlying handlebody $H$ embeds into $M$ in a nice way. Any such orbifold cover is defined by a finite index subgroup $\Gamma\leq\pi_{1}^{\operatorname{orb}}\left(H\right)$, and inherits a hyperbolic structure by lifting. We want to find one that satisfies certain geometric conditions on the lifted hyperbolic structure, needed to apply the results from [KM] about the existence of negatively curved metrics on the cone-off, see §2.1. Specifically, we must bound on the injectivity radius of the boundary $\partial M$ from below. (We warn the reader that when taking an orbifold cover $M\to H$, the systole of $\partial M$ could be smaller than the systole of $\partial H$ as defined in Definition 4.2.) To achieve our lower bound, we will exploit the residual finiteness of $W$ to carefully choose a finite-index torsion-free subgroup $K$ of $W$ and let $M$ be the orbifold cover of $H$ corresponding to $\Gamma=f^{-1}(K)$, where the Coxeter group $W$ and the map $f:\pi_{1}^{\operatorname{orb}}\left(H\right)\to W$ were defined in §4.2.
@@
Since $\operatorname{w}(\gamma)$ is not reduced, it is not the empty word and there is a subword of the form $t\omega t$, where $\omega$ is a (possibly empty) word in the only two generators $t_{i},t_{j}\neq t$ that commute with $t$, see [DA08, §3.4]. (Note that each topological boundary component of $S$ consists of at least $6$ mirrors, by Assumption 4.4 and Lemma 2.8.) This means that there is a subarc $\alpha$ of $\gamma$ which starts and ends on the mirror $t$ so that any sufficiently small geodesic extension $\alpha_{\epsilon}$ of $\alpha$ satisfies $\operatorname{w}(\alpha_{\epsilon})=t\omega t$. (Here $\alpha_{\epsilon}$ is the geodesic arc obtained by extending $\alpha$ geodesically at each end by $\epsilon/2$.)
@@
Since $\operatorname{w}(\gamma)$ is reduced, for any subword of $\operatorname{w}(\gamma)$ of the form $t_{i}t_{j}t_{k}$ we have that either $i,j,k$ are all distinct, or $t_{i}=t_{k}$ does not commute with $t_{j}$, see [DA08, §3.4]. In the first case, at least two of the mirrors are disjoint, so we get a contribution of $D$ to the length of $\gamma$. In the second case $t_{j}$ is not adjacent to $t_{i}=t_{k}$, hence we get contribution of $D$ in this case as well. Since we get a contribution of at least $D$ to the length of $\gamma$ from each subword of length $3$ of $\operatorname{w}(\gamma)$, it follows that $\ell(\gamma)\geq\left\lfloor{\frac{|\operatorname{w}(\gamma)|_{T}}{3}}\right\rfloor D$. ∎

### v2

@@ Abstract.
We construct compact hyperbolic $3$-manifolds with totally geodesic boundary, such that the closed $3$-pseudomanifolds obtained by coning off the boundary components are negatively curved and contain locally convex subspaces whose fundamental groups have property (T). In particular, the fundamental groups of these $3$-pseudomanifolds are word hyperbolic but not cubulable. We deduce that in any relative cubulation of one of these hyperbolic $3$-manifold groups some hyperplane stabilizer has infinite intersection with the fundamental group of some boundary component.
@@ 1. Introduction
To say a group is cubulable is to say it admits a proper and cocompact action by isometries on some $\textrm{CAT}(0)$ cube complex. Every cocompact Kleinian group is cubulable by [5, Theorem 1.5]. And a cubulable word hyperbolic group with $2$-sphere boundary is virtually cocompact Kleinian by [28, Theorem 1.10] (generalizing [41]).
We do not directly speak to Question 1.1, but we study the corresponding question for groups with Pontryagin sphere boundary. The Pontryagin sphere is a certain inverse limit of surfaces which appears as the boundary at infinity of many $\textrm{CAT}(0)$ 33-dimensional complexes (universal covers of $3$-pseudomanifolds, see below). It is like a sphere in that it is a connected $2$-dimensional compactum whose $2$-dimensional Čech cohomology is cyclic. However its $1$-dimensional cohomology is infinitely generated. We refer the reader to [29, 18, 24, 47].
@@
Classical examples of word hyperbolic groups with Pontryagin sphere boundary occur for instance as finite index subgroups of a right-angled Coxeter group (RACG) whose nerve is a flag-no-square triangulation of a connected closed orientable surface, and via Charney–Davis strict hyperbolization [16, 11]. Both sources of examples yield groups that are cubulable (see [16, 34]). Nonetheless, we answer Question 1.2 in the negative; see (1) in Corollary 1.4.
@@
The existence of an infinite subgroup with property (T) is a well-known obstruction to the existence of nice geometric actions, and makes these groups quite different from $3$-manifold groups. More precisely, we have the following. (See [4] for the definitions of the Haagerup and (T) properties, and [1] for the definition of RFRS.)
@@
The non-cubulability of $\pi_{1}(\widehat{M}_{n})$ does not automatically imply the non-cubulability of fundamental groups of $3$-pseudomanifolds obtained by coning off the boundary components of an arbitrary finite-sheeted cover of $M_{n}$. Indeed, one may use Wise’s Malnormal Special Quotient Theorem [48, Theorem 12.2] together with the cubulability of $\pi_{1}(M_{n})$ [48, Theorem 17.14] to see that each $M_{n}$ has some finite sheeted cover $M_{n}^{\prime}$ so that $\pi_{1}(\widehat{M}_{n}^{\prime})$ is cubulable.
@@
The Gromov boundary of the groups $\pi_{1}(\widehat{M}_{n})$ in Theorem 1.3 is the tree of manifolds defined by closed orientable surfaces of positive genus, i.e., a Pontryagin sphere, see [33, Theorem A.1] and [47]. Moreover, the Gromov boundary of the infinite quasiconvex subgroup with property (T) has limit set a Menger curve, see [31].
@@
The groups $\pi_{1}(\widehat{M}_{n})$ in Theorem 1.3 split over quasiconvex surface subgroups (see Proposition 5.1), so they do not have property (T). In particular, they are $3$-dimensional word hyperbolic groups without (T) and without Haagerup, whose boundary is connected and has cyclic top-dimensional Čech cohomology (since the boundary is a Pontryagin sphere, see [47, Theorem 6.2]). See [35, Remark 5.6] for other examples in dimension $\geq 9$, whose boundary is a topological sphere. For examples of non-Kleinian hyperbolic 3-pseudomanifold groups that have the Haagerup property, and even act convex cocompactly on $\mathbb{H}^{n}$ for some $n>3$, see the RACGs considered in [17].
Remark 1.8 (Relatively geometric cubulation).
Ian Agol pointed out the following consequence of our construction (see Proposition 5.2): For any relatively geometric cubulation (in the sense of [20]) of any of the Kleinian groups $\pi_{1}(M_{n})$ that we construct, there must be a hyperplane stabilizer with infinite intersection with some boundary subgroup of $\pi_{1}(M_{n})$. This is in contrast to the situation for finite volume cusped 33–manifolds. The fundamental groups of such manifolds admit relatively geometric cubulations by quasi-Fuchsian closed surface subgroups by the ubiquity results in [12, 30].
@@
In §2 we fix our notation and terminology about the construction of pseudomanifolds as cone-offs of manifolds and about orbifolds. Moreover, we collect some preliminary material from [37] about the $2$-dimensional simplicial complexes $T_{n}$ whose fundamental groups have property (T). In §3 we present the construction of a particular hyperbolic $3$-orbifold with boundary, and in §4 we show how to construct suitable finite covers for which we can obtain quantitative control on various geometric quantities. The orbifold structure constructed in §3 is given by orthogonal mirrors. This is reminiscent of the Davis reflection trick [15]. However, we use the mirror structure to construct manifolds with boundary instead of closed manifolds. The geometric control enables us to use the results from [33] to construct negatively curved metrics on the cone-offs of such covers. The proofs of the main theorem and corollary stated in the Introduction are presented in §5. We end the paper with some open questions in §6.
Acknowledgments. J.M. was partially supported by the Simons Foundation, grant #942496. L.R. was partially supported by INDAM-GNSAGA. We thank Chris Hruska for pointing us to useful references. We thank Ian Agol for pointing out Proposition 5.2, and Daniel Groves for useful conversations.
@@ 2.1. Coning off
Here we state the results from [33] which allow us to put negatively curved (i.e., locally $\textrm{CAT}(k)$ for some $k<0$) metrics on our $3$-pseudomanifolds. Let $M$ be a compact $3$-manifold with boundary. The cone-off is the space
@@
Sufficient geometric conditions are given in [33] for $\widehat{M}$ to admit a negatively curved metric $\hat{d}$ and for the induced cone-off of a subspace $Z$ to be locally convex. To state the conditions we need a definition from [33]:
@@
[33, Theorem A] Let $M$ be a compact hyperbolic manifold with totally geodesic boundary, let $b$ be a positive number less than $\operatorname{BW}_{M}(\partial M)$ and let $c>\pi/\sinh(b)$. Suppose
@@
The following is a special case of a result proved in [33]:
@@
[33, Theorem B] Let $M$, $b$, and $c$ satisfy the hypotheses of Theorem 2.2 and let $\hat{d}$ be the negatively curved metric on $\widehat{M}$ constructed in that theorem. Suppose $Z\subset M$ is closed, locally convex, is a tame product near $N_{b}(\partial M)$, and that
@@ 2.2. Orbifolds
For an introduction to the general notion of a smooth orbifold see [7, 9, 6, 13, 32]. We briefly recall the idea: An $n$-orbifold (possibly with boundary) consists of a topological space (the underlying space) together with an open cover by sets of the form $U=\tilde{U}/\Gamma$, where $\tilde{U}$ is an open subset of $\mathbb{R}^{n}$ or $\mathbb{R}^{n-1}\times[0,\infty)$, and $\Gamma$ is a finite group of diffeomorphisms with nonempty fixed point set. (Smoothness also implies that a nonempty intersection between these sets gives rise to smooth transition maps.) If a point $x$ in the underlying space is the image of a fixed point of $\Gamma$, we say that the isotropy group of $x$ is $\Gamma$. (Strictly speaking we replace $\Gamma$ by the finite subgroup of $O(n)$ given by taking the derivative of its elements; this gives a well-defined group up to conjugacy in $O(n)$.) The boundary $\partial X$ is the (n1)(n-1)-orbifold whose underlying space consists of those points where the orbifold chart always comes from $\mathbb{R}^{n-1}\times[0,\infty)$, and whose isotropy groups are the isotropy groups for $X$, restricted to $\mathbb{R}^{n-1}\times\{0\}$. The isotropy groups of the points in the boundary lie in O(n1)O(n-1). The orbifold $X$ is said to be closed if its underlying space is compact and $\partial X$ is empty.
@@ 2.3. The seed with property (T)
We describe here a construction which is closely related to the examples in [37]. In particular we will describe some fundamental groups of triangle complexes which are subgroups of finite index in some of the groups considered there.
@@
Let $\Theta$ be the $1$-complex with two vertices and $k$ edges between them. The following is a consequence of work of Lubotzky–Phillips–Sarnak [36] and Margulis [40].
@@
This is explained in [37], following closely [38, Chapter 7]. For the fact that the graphs can be taken in a tower, see the bottom of page 64 of [37]. The reason is that the subgroups of $\pi_{1}(\Theta)$ corresponding to the covers $\Lambda_{n}\to\Theta$ arise as congruence subgroups $\Gamma(2q_{n})$ for an arithmetic group $\Gamma\cong\pi_{1}(\Theta)$, and the numbers $q_{n}$ can be taken to be successive powers of a prime $q$ where k1k-1 is not a square modulo $q$. (For example if k1=17k-1=17 we could take $q_{n}=3^{n}$ or $q_{n}=5^{n}$.) ∎
@@
Since $\tau$ is induced by a graph automorphism of $\Theta$, all the vertex links in $\widehat{T}_{n}$ are isomorphic as graphs to the cover $\Lambda_{n}$. By Theorem 2.6, for $n$ large enough we can assume that the link of every vertex in $\widehat{T}_{n}$ has girth at least $6$. Now, if we identify each triangle of $\widehat{T}_{n}$ with an equilateral Euclidean triangle, then the induced metric is locally $\textrm{CAT}(0)$, so $\widehat{T}_{n}$ is aspherical. This asphericity, together with the condition on the Laplacian of $\Lambda_{n}$ in Theorem 2.6, allows us to apply [3, Corollary 1] (or [49, Théorème 1]) to deduce that $\pi_{1}(\widehat{T}_{n})$ has property (T). ∎
@@ 3. The handlebody $H_{0}$ and its hyperbolic orbifold structure
To recall the strategy of our main construction: We will first take the cover $p:H_{n}\to H_{0}$ corresponding to a cover $T_{n}\to T_{0}$ from Definition 2.7, with $n$ large enough to apply Lemma 2.8. We will then take an orbifold cover $M$ of $H=H_{n}$ to which $H=H_{n}$ lifts homeomorphically, and then cone off $\partial M$. Choosing these covers carefully, and using the results from [33] (see §2.1) the inclusion of the induced cone-off $\widehat{H}$ into $\widehat{M}$ will have $\pi_{1}$-image isomorphic to $\pi_{1}(\widehat{T}_{n})$, which has property (T) by Lemma 2.8.
@@ 3.1. The hyperbolic orbifold structure on $H_{0}$
The underlying space of $B$ is a closed ball, but the orbifold structure suggests an abstract polyhedron, with dihedral angles of $\pi/2$ at all edges involving the unmirrored face $F_{B}$, and $\pi/k$ at each edge with isotropy group dihedral of order $2k$. One can see by direct inspection that this polyhedron satisfies the hypotheses of Andreev’s Theorem (see [46, Theorem 1.4], which corrects some mistakes in the original [2]). In particular, there is a unique hyperbolic polyhedron with the specified dihedral angles, and we identify this polyhedron with $B$. Mirroring all the faces with an acute adjacent dihedral angle gives a hyperbolic orbifold with the same orbifold structure specified above on $B$. In other words the polyhedron determines a hyperbolic structure on the orbifold $B$. Since the dihedral angles adjacent to the boundary face $F_{B}$ are $\pi/2$, this structure has totally geodesic boundary.
@@ 4.1. The orbifold $H=H_{n}$ covering $H_{0}$
Each component of $\partial H_{n}$ is a $2$-dimensional thickening of a cover of a theta graph. A cover of a theta-graph can be non-planar (eg $K_{3,3}$), so it is possible that these components have positive genus. Indeed, they must: if they were planar, then the coned-off complex would be homotopy equivalent to a $3$-manifold with boundary, so its fundamental group could not contain an infinite subgroup with property (T), see [25].
@@ 4.3. The manifold $M$ orbifold covering $H$
We want to construct a manifold $M$ (with empty orbifold locus) that is an orbifold cover of the orbifold $H$, so that the underlying handlebody $H$ embeds into $M$ in a nice way. Any such orbifold cover is defined by a finite index subgroup $\Gamma\leq\pi_{1}^{\operatorname{orb}}\left(H\right)$, and inherits a hyperbolic structure by lifting. We want to find one that satisfies certain geometric conditions on the lifted hyperbolic structure, needed to apply the results from [33] about the existence of negatively curved metrics on the cone-off, see §2.1. Specifically, we must bound on the injectivity radius of the boundary $\partial M$ from below. (We warn the reader that when taking an orbifold cover $M\to H$, the systole of $\partial M$ could be smaller than the systole of $\partial H$ as defined in Definition 4.2.) To achieve our lower bound, we will exploit the residual finiteness of $W$ to carefully choose a finite-index torsion-free subgroup $K$ of $W$ and let $M$ be the orbifold cover of $H$ corresponding to $\Gamma=f^{-1}(K)$, where the Coxeter group $W$ and the map $f:\pi_{1}^{\operatorname{orb}}\left(H\right)\to W$ were defined in §4.2.
@@
Since $\operatorname{w}(\gamma)$ is not reduced, it is not the empty word and there is a subword of the form $t\omega t$, where $\omega$ is a (possibly empty) word in the only two generators $t_{i},t_{j}\neq t$ that commute with $t$, see [16, §3.4]. (Note that each topological boundary component of $S$ consists of at least $6$ mirrors, by Assumption 4.4 and Lemma 2.8.) This means that there is a subarc $\alpha$ of $\gamma$ which starts and ends on the mirror $t$ so that any sufficiently small geodesic extension $\alpha_{\epsilon}$ of $\alpha$ satisfies $\operatorname{w}(\alpha_{\epsilon})=t\omega t$. (Here $\alpha_{\epsilon}$ is the geodesic arc obtained by extending $\alpha$ geodesically at each end by $\epsilon/2$.)
@@
Since $\operatorname{w}(\gamma)$ is reduced, for any subword of $\operatorname{w}(\gamma)$ of the form $t_{i}t_{j}t_{k}$ we have that either $i,j,k$ are all distinct, or $t_{i}=t_{k}$ does not commute with $t_{j}$, see [16, §3.4]. In the first case, at least two of the mirrors are disjoint, so we get a contribution of $D$ to the length of $\gamma$. In the second case $t_{j}$ is not adjacent to $t_{i}=t_{k}$, hence we get contribution of $D$ in this case as well. Since we get a contribution of at least $D$ to the length of $\gamma$ from each subword of length $3$ of $\operatorname{w}(\gamma)$, it follows that $\ell(\gamma)\geq\left\lfloor{\frac{|\operatorname{w}(\gamma)|_{T}}{3}}\right\rfloor D$. ∎